# Analysis of Broken Randomized Experiments by Principal Stratification

### v1

@@ 5 Application to Job Corps data
Year Employment $S$ Earnings $Y$
@@
Figure 1 shows the estimated treatment effect on employment $E\{S(1)-S(0)\mid c\}$ and the estimated treatment effect on weekly earnings $E\{Y(1)-Y(0)\mid cl\}$ with 95% confidence intervals. In the first year after assignment, there is a negative effect on employment. This is possibly because the individuals spent their time on training and they were not ready to enter the job market in the first year. Since the third year after assignment, there is a positive effect on employment, indicating that the training program helped the youths finding jobs. However, the effect is not significant at level 0.05. For the always-employed people, the training program has a significant effect on earnings.
@@
Table 4: The estimated mean outcomes, treatment effect, standard error (se) and $P$-value in different years by the proposed method (PACE), two-stage least square (2SLS), intention-to-treat analysis (ITT), as-treated analysis (AT) and per-protocol analysis (PP) for survived compliers
Year Method $\widehat{\mu}_{1}$ se $\widehat{\mu}_{0}$ se $\widehat{\tau}$ se $P$
1 PACE 4.902 (0.090) 5.008 (0.043) -0.106 (0.101) 0.291
2SLS -0.128 (0.080) 0.107
ITT 4.859 (0.018) 4.903 (0.020) -0.044 (0.027) 0.105
AT 4.800 (0.016) 5.040 (0.023) -0.240 (0.028) 0.000
PP 4.813 (0.020) 5.027 (0.027) -0.214 (0.034) 0.000
2 PACE 4.855 (0.082) 4.879 (0.050) -0.025 (0.096) 0.798
2SLS -0.026 (0.095) 0.788
ITT 4.737 (0.017) 4.744 (0.021) -0.007 (0.027) 0.789
AT 4.704 (0.015) 4.860 (0.027) -0.156 (0.031) 0.000
PP 4.723 (0.018) 4.868 (0.033) -0.145 (0.037) 0.000
3 PACE 5.211 (0.068) 5.027 (0.046) 0.185 (0.081) 0.022
2SLS 0.207 (0.082) 0.012
ITT 5.016 (0.014) 4.958 (0.019) 0.058 (0.023) 0.013
AT 4.993 (0.013) 4.996 (0.025) -0.003 (0.028) 0.921
PP 5.023 (0.015) 5.009 (0.030) 0.013 (0.034) 0.687
4 PACE 5.305 (0.067) 5.185 (0.047) 0.120 (0.081) 0.137
2SLS 0.127 (0.082) 0.121
ITT 5.199 (0.014) 5.163 (0.018) 0.036 (0.023) 0.121
AT 5.193 (0.013) 5.156 (0.025) 0.037 (0.028) 0.184
PP 5.207 (0.015) 5.168 (0.030) 0.039 (0.033) 0.239
Table 4 summarizes the estimated mean outcomes, treatment effect on weekly earnings, standard error (se) and $P$-value in different years by the proposed method (PACE), two-stage least square (2SLS), intention-to-treat analysis (ITT), as-treated analysis (AT) and per-protocol analysis (PP) for survived compliers. According to the proposed estimator, in the third year after assignment, there is a significant and positive effect on weekly earnings. The always-employed and compliers can have 18.5% higher salaries if taking the training program. The effect slightly goes smaller in the fourth year because the weekly earnings without participating in the training program $\mu_{0}$ got higher. Different methods give almost the same signs of the treatment effect. In the third and forth year, ITT, AT and PP underestimate the principal stratum treatment effect, indicating that compliers can benefit more from the training program. The proposed method yields the largest standard error among all the methods under comparison, because we define the estimand in a specific principal stratum, which has the fewest individuals among all methods. The point estimate of 2SLS is similar to that of PACE, indicating that the cross-world reliance of employment on the compliance type is weak (which is consistent with the left panel of Figure 1). The 2SLS has slightly smaller variance than PACE.
We can further incorporate covariates in the estimation. We choose six baseline covariates: gender, age, log household size, years of education, ever worked and health. There are missing values in covariates. We perform 10 times of imputation. In each time, we use simple imuptation to fill the missinge values by drawing values from the observed covariates distribution. Then we use linear regression to model the earnings $Y$ on covariates. Finally, we use the Rubin’s rule to calculate the sampling variance. Figure 2 shows the stimated treatment effect on earnings. Compared with Figure 1, we obtain tighter confidence intervals. The treatment effects are significant in the third and fourth years. The $P$-values are 0.0152, 0.6100, 0.0000 and 0.0036 for these four years, respectively. It is interesting that the treatment effect is significantly negative in the first year, indicating that joining the training program leads to lower earnings, probably because that the individuals in the program have shorter work time and hence less earnings. Figure 3 presents the estimated treatment effects by other methods with covariates adjustment. The trends are similar but the magnitudes differ. The confidence intervals given by 2SLS, ITT, AT and PP are tighter because these four methods do not account for the uncertainty in estimating the probability of survived (always-employed) compliers.
Figure 2: Time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$, with covariates adjustment.
Figure 3: Time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$ estimated by 2SLS, ITT, AT and PP, with covariates adjustment.
@@
The results by the proposed method have meaning causal interpretations. However, since the target population, consisting of the survived compliers, is not individually identifiable, how to guide the results for future policy making is doubtful. Although we know the proportion of survived compliers, we do not know who will benefit from the training program. From the perspective of policy makers, the intention-to-treat analysis is the easiest to interpret, as it basically evaluates the effect of the treatment policy, regardless of whether the individuals adhere to the treatment assignment. In contrast, the proposed method evaluates a pure effect of the training program. The estimate by the proposed method is only valid for survived compliers, so we can never know how things would be if the training is made compulsory. Nevertheless, different methods approaches the scientific question from different perspectives, and assessment by different methods provides a comprehensive knowledge of how the treatment can make an effect.
@@ 6 Discussion
Post-treatment events bring challenges to conduct causal analysis of the treatment effect. Although the baseline characteristics are naturally balanced by randomization, these characteristics are not balanced anymore if there is post-treatment selection. In this paper, we consider the broken randomized experiments with non-compliance, truncation-by-death and missing data. Non-compliance leads the received treatment different from the assigned treatment, and the outcomes after the received treatment are still meaningful, reflecting a mixed effect of the treatment and possibly unmeasured compliance type. Truncation-by-death renders the outcomes ill defined or undefined. The survival type also embeds information about underlying characteristics of individuals. When there are non-compliance, truncation-by-death and missingness simultaneously in a randomized trial, we propose a framework to identify and estimate the principal stratum average treatment effect. The causal effect is defined in a subpopulation called the survived compliers. Although we can neither identify who belongs to this subpopulation or identify the proportion of this subpopulation, we are able to identify the treatment effect in this subpopulation.
@@
There are some limitations for the estimand. First, the identification involves untestable assumptions. The first four assumptions on non-compliance are in line with the assumptions to identify the complier average causal effect (CACE) or the local average treatment effect (LATE). Lots of efforts have been devoted to find testable implications for these assumptions (Yang et al., , 2014; Mogstad et al., , 2018; Kédagni and Mourifié, , 2020). The principal ignorability assumption is hard to justify, and sensitivity analysis may be helpful to assess the robustness of substantial results to violation of principal ignorability. Utilizing additional information such as substitutional variables or post-treatment correlates, principal ignorability can be relaxed (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
@@
Second, we did not consider adjustments of measured covariates, especially in the identification assumptions. More efficient estimator can be proposed by incorporating covariates in the modeling. However, relaxing ignorability assumptions by conditioning on covariates could incur other challenges because the average potential outcomes need to be integrated with respect to the distribution of covariates in the target principal stratum while the proportion of this principal stratum given covariates has not been identified. In Appendix, we provide some identification results with baseline covariates. This issue is important if we aim to extend the proposed framework to observational studies.
@@
Third, as a common criticism to principal stratification, it is hard to generalize the result to a new population. Even if the policy is effective, policy makers cannot easily implement such a policy because we never know who will follow the policy instructions and benefit from the policy. As a tool, principal stratification defines the estimand in a subpopulation where all potential outcomes are well defined and comparabe. As a goal, the principal stratum shuold be consistent with the scientific interest. In order that the estimated treatment effect has a meaningful interpretation, the application scenario should be careful justified.
@@
The data we use in our work is publicly available in the R package causalweight. The R codes that support our findings are available as part of the online supplementary material.
Acknowledgements
We thank the support from the Save 2050 Programme jointly sponsored by Swarma Club and X-Order. JEL: C18, C39, C51
Funding information
This work is supported by the National Key Research and Development Program of China, Grant No. 2021YFF0901400; National Social Science Fund of China, Grant No. 19BJY108.
Appendix: Identification of the causal effect with covariates
For simplicity, suppose there are no missing values, otherwise multiple imputation approaches can be applied before data analysis. Here we restate the identification assumptions. The independencies in Assumptions 16 should be conditioned additionally on baseline covariates $X$, as listed below.
For notation simplicity, we assume there is no missingness (otherwise we just need to add the missingness indicator in the conditional expectations). In the presence of covariates, one can easily identify and estimate the conditional principal average causal effect (or called the local covariate-specific average treatment effect),
$E\{Y(1)\mid G=cl,X\}$ $=\frac{E\{DSY\mid Z=1,X\}-E\{DSY\mid Z=0,X\}}{E\{DS\mid Z=1,X\}-E\{DS\mid Z=0,X\}},$
$E\{Y(0)\mid G=cl,X\}$ $=\frac{E\{(1-D)SY\mid Z=1,X\}-E\{(1-D)SY\mid Z=0,X\}}{E\{(1-D)S\mid Z=1,X\}-E\{(1-D)S\mid Z=0,X\}}.$
However, when it comes to the population-level average causal effect, the main obstacle for identification is that the proportion of the $cl$ stratum is not identifiable. Indeed, solely with principal ignorability, one can never distinguish $cl$ and $cp$ strata among the treated, nor distinguish $cl$ and $ch$ strata among the controlled using observed data. Therefore, additional assumptions should be made to identify the population-level estimand.
Assumption 7 (Monotonicity on survival)
$S(1)\geq S(0)$ almost surely.
Under monotonicity on survival, the proportions of each principal stratum are all identifiable. Figure 4 illustrates the constitution of observed data in each group. Particularly, the proportion of the $cl$ stratum conditional on covariates is identified as
$P(G=cl\mid X=x)$ $=P(D=0\mid Z=0)E\{S\mid(1-Z)(1-D)\Delta_{S}=1,X=x\}$
$\qquad-P(D=0\mid Z=1)E\{S\mid Z(1-D)\Delta_{S}=1,X=x\}.$
The first term is the conditional proportion of $cl$ and $nl$, and the second term is the conditional proportion of $nl$. Next, the population-level principal average causal effect can be obtained by taking expectation for the conditional principal average causal effect over the covariate distribution on the $cl$ stratum.
Figure 4: Monotonicity on survival. Dark shaded are the observed survivors.
Theorem 4
Under the conditional version of Assumptions 16 and Assumption 7, the principal average causal effect $\tau$ is nonparametrically identifiable,
[…]
where $P(x)$ is the distribution function of $X$ and $\mathcal{X}$ is the support of $X$.
The identification formula motivates a regression estimator for the target estimand. In fact, one can construct weighting estimators using the propensity score and principal score (Feller et al., , 2017; Ding and Lu, , 2017). Then multiply robust estimator can be derived by finding the efficient influence function of the estimand (Wang and Tchetgen Tchetgen, , 2018; Jiang et al., , 2022). When models are correctly specified, the efficient influence function based estimator may enjoy better efficiency than the regression estimator and weighting estimator.
Monotonicity is untestable. Whether monotonicity on survival is reasonable depends on specific scientific questions. There are some variants of the individual-level monotonicity such as stochastic monotonicity (Lee et al., , 2010), which assumes a known relation, also relying untestable prior knowledge, between the potential survivals. To relax the principal ignorability assumption, we can invoke a substitutional (negative control) variable related with potential survivals but not directly related with potential outcomes (Wang et al., , 2017; Deng et al., , 2021). The substitutional variable embeds information of the principal stratum, so the different distribution of mean outcomes can be separated by recovering the information of $G$ using $V$. The data structure would be more complex and may not be easily interpretable for actual scenarios.

### v2

@@ 5 Application to Job Corps study
Year Employment $S$ Earnings $Y$ in logarithm
@@
We control six baseline covariates to model the earnings: gender, age, log household size, years of education, ever worked, and health. There are missing values in covariates. We perform 10 times of imputation. Finally, we use Rubin’s rule to calculate the sampling variance. Figure 2 shows the estimated treatment effect on employment $E\{S(1)-S(0)\mid c\}$ and the estimated treatment effect on weekly earnings $E\{Y(1)-Y(0)\mid cl\}$ with 95% confidence intervals. In the first year after the assignment, there is a negative effect on employment. This is probably because the individuals spent their time on training, and they were not ready to enter the job market in the first year. Since the third year after the assignment, there has been a positive effect on employment, indicating that the training program helped the youths find jobs. The $P$-values are 0.0000, 0.8813, 0.0079, and 0.0030 for the four years, respectively.
@@
The treatment effect on earnings is negative in the first year, indicating that joining the training program leads to lower earnings, probably because the individuals in the program have shorter work time. In the third year after the assignment, there is a significant and positive effect on earnings. The always-employed and compliers have 18.5% higher salaries if they take the training program. The effect becomes slightly smaller in the fourth year because the weekly earnings without participating in the training program $\mu_{0}$ got higher. The $P$-values are 0.2700, 0.7885, 0.0140, and 0.1224 for these four years, respectively.
Figure 3 presents the estimated treatment effects by other methods. The point estimate of 2SLS is similar to that of the proposed method, indicating that the cross-world reliance of employment on the compliance type is weak. In the third and fourth years, ITT, AT, and PP underestimate the principal stratum treatment effect, indicating that compliers can benefit more from the training program. The confidence intervals given by ITT, AT, and PP are tighter because these methods do not account for the uncertainty in estimating the probability of survived (always-employed) compliers.
Figure 3: Time-varying treatment effect on earnings $E\{Y(1)-Y(0)\mid cl\}$ estimated by 2SLS, ITT, AT and PP.
@@
The results of the proposed method have meaningful causal interpretations. However, since the target population, consisting of the survived compliers, is not individually identifiable, how to guide the results for future policy making is doubtful. Although we know the proportion of survived compliers, we do not know who will benefit from the training program. From the perspective of policymakers, the intention-to-treat analysis is the easiest to interpret, as it basically evaluates the effect of the treatment policy, regardless of whether the individuals adhere to the treatment assignment. In contrast, the proposed method evaluates the pure effect of the training program. The estimate by the proposed method is only valid for survived compliers, so we can never know how things would be if the training is made compulsory. Nevertheless, different methods approach the scientific question from different perspectives, and assessment by different methods provides a comprehensive knowledge of how the treatment can make an effect. In the Supplementary Material, we relax the identification assumptions by adjusting for baseline covariates and infer an interventionist complier average treatment effect. We find that the treatment effect on earnings is significant for the third and fourth years in compliers.
@@ 6 Discussion
Post-treatment events bring challenges to conducting causal analyses of the treatment effect. Although the baseline characteristics are naturally balanced by randomization, these characteristics are not balanced anymore if there is post-treatment selection. In this paper, we consider the broken randomized experiments with non-compliance, truncation-by-death, and missing data. Non-compliance leads the received treatment to differ from the assigned treatment, and the outcomes after the received treatment are still meaningful, reflecting a mixed effect of the treatment and possibly unmeasured compliance type. Truncation-by-death renders the outcomes ill-defined or undefined. The survival type also embeds information about the underlying characteristics of individuals. When there are non-compliance, truncation-by-death, and missingness simultaneously in a randomized trial, we propose a framework to identify and estimate the principal stratum average treatment effect. The causal effect is defined in a subpopulation called the survived compliers. Although we cannot identify who belongs to this subpopulation or the proportion of this subpopulation, we can still identify the treatment effect in this subpopulation.
@@
There are some limitations for the estimand. First, the identification involves untestable assumptions. The four assumptions on non-compliance are in line with the assumptions to identify the complier average causal effect (CACE) or the local average treatment effect (LATE). Many efforts have been devoted to finding testable implications for these assumptions (Yang et al., , 2014; Mogstad et al., , 2018; Kédagni and Mourifié, , 2020). The principal ignorability assumption is hard to justify, and sensitivity analysis may be helpful to assess the robustness of substantial results to the violation of principal ignorability. A conservative approach is to find bounds for the treatment effect (Long and Hudgens, , 2013; Marden et al., , 2018). Whether the bounds are informative is worthy of discussion. Utilizing additional information such as substitutional variables or post-treatment correlates, principal ignorability can be relaxed (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
@@
Second, we did not consider adjustments of measured covariates, especially in the identification assumptions. Covariate adjustment allows for studying causal effects in observational studies and fosters more efficient estimation (Feller et al., , 2017; Jiang et al., , 2022). However, relaxing ignorability assumptions by conditioning on covariates could incur other challenges because the average potential outcomes need to be integrated over the distribution of covariates in the target principal stratum, while the proportion of this principal stratum given covariates has not been identified. In Supplementary Material B, we provide some identification results with baseline covariates, under an additional survival monotonicity assumption. This issue is crucial if we aim to extend the proposed framework to observational studies. In Supplementary Material C, we propose an interventionist estimand on compliers, which does not need survival monotonicity.
@@
Third, as a common criticism of principal stratification, it is hard to interpret the causal effect and generalize the result to a new population (Pearl, , 2011; VanderWeele, , 2011). The causal effect is defined on a latent subpopulation where both potential outcomes theoretically exist. An individual in this subpopulation must comply with treatment assignments and be destined to have an outcome. Even if a policy is effective, policymakers cannot easily implement it because they never know who will follow the policy instructions and benefit from the policy. As a tool, principal stratification defines the estimand in a subpopulation where all potential outcomes are well-defined and comparable. As a goal, the principal stratum should be consistent with the scientific interest. The application scenario should be carefully justified so that the estimated treatment effect can be meaningfully interpreted.
Acknowledgements
We thank the associate editor and reviewers for theor comments. We thank the support from the Save 2050 Programme jointly sponsored by the Swarma Club and X-Order. JEL: C18, C39, C51
Funding information
This work is supported by the National Social Science Fund of China, Grant No. 19BJY108.
@@ Acknowledgements
The data we use in our work is publicly available in the R package causalweight.
Supplementary Material
The supplementary material includes estimands for binary outcomes and estimands under assumptions adjusting for baseline covariates.